OpenStudy (sierra_grey):

im not sure how to do this. Two consecutive integers have a sum of 157. Find the integers.

OpenStudy (anonymous):

Let x equal the first integer. So the next integer as it relates to x would be what?

OpenStudy (anonymous):

x = first integer x+1 = second integer (since its consecutive) Both integers equal 157 so.. x + x + 1 = 157

OpenStudy (anonymous):

Are there any like terms that you can combine?

OpenStudy (anonymous):

x + x = ?

OpenStudy (sierra_grey):

first integer = x = 78 second integer = x+1 = 78+1 = 79 got it. thanks everyone!

OpenStudy (anonymous):

Good job!